Cutting a stick A stick n inches long needs to be cut into n 1-inch pieces. Outline an algorithm that performs this task with the minimum number of cuts if several pieces of the stick can be cut at the same time. Also give a formula for the minimum number of cuts
Answer #1

the same time. the approach soly since we can cut several pieces at to solve this problem: cut the stick into 2 pleces : OneAlgorithm:- c= 0 8=n / intial value for r . S=O sls minimum no. ofcuts c=int (10qan) s=sto &=n-2c if (r=0) printNo. of cuts